# Neural Green’s Operators for Parametric Partial Differential Equations

### v4

@@ 5.2 Advection-Diffusion
The final test case we consider is the advection-diffusion equation with inhomogeneous Dirichlet and Neumann boundary conditions:
@@
where $\left(\mathbf{c}\cdot\mathbf{n}\right)_{-}=\min\left(\mathbf{c}\cdot\mathbf{n},0\right)$. Compared to the pure diffusion equation, the advection-diffusion equation presents a more challenging problem due to a number of factors. First, the solution now depends non-linearly on both the diffusion $\theta$ and the velocity field $\mathbf{c}$. Second, for large Péclet numbers (the ratio between the advection $\mathbf{c}$ and the diffusion $\theta$) the true solution will feature boundary layers near the Dirichlet boundaries, which may be difficult to resolve. Note that the boundary condition $\theta\nabla u\cdot\mathbf{n}-\left(\mathbf{c}\cdot\mathbf{n}\right)_{-}u=\eta$ differs from the Neumann boundary condition on the inflow part of the top and bottom boundaries (i.e. where $\mathbf{c}\cdot\mathbf{n}<0$). This change was made to facilitate stable finite element discretizations.
@@ 5.2.2 Results
Figure 20(b) shows the true solution alongside the solution predictions made by the different model architectures for one problem from the training data, and one from the testing data.
@@
Figure 20: The true solution and model predictions to one problem from the training data set (Figure 20(a)) and one problem from the testing data set (Figure 20(b))
@@
To illustrate the scaling of the error with the data-set size we show in Figure 21 the effect that the amount of training data has on the model accuracy for FNOs, CNOs, data NGOs, and model NGOs. As the plot shows, both NGO models outperform the other models for each tested data set size, although the difference becomes small for large data sets.
@@ 6 Conclusion
In this work, we have introduced the Neural Green’s Operator, a machine learning architecture for learning the solution operator to parametric families of linear PDEs. Our construction of NGOs is based on a finite-dimensional representation of Green’s functions. Such a representation enables the preservation of the linear action of Green’s operators on the inhomogeneity fields, while approximating the nonlinear dependence of the Green’s function on the coefficients of the PDE using neural networks that take weighted averages of such coefficients as input. This construction reduces the complexity of the problem from learning the entire solution operator and its dependence on all parameters to only learning the Green’s function and its dependence on the PDE coefficients. Moreover, taking weighted averages, rather than point samples, of input functions decouples the network size from the number of sampling points, enabling efficient resolution of multiple scales in the input fields. We distinguished three use cases of NGOs: the model NGO, which is used when the underlying PDE is known and training data is available; the data-free NGO, which is used when the underlying PDE is known, but training data is unavailable; and the data NGO, which is used when training data is available, but the model that underlies the data is unknown. For the steady diffusion problem, the accuracy of NGOs exceeds that of VarMiONs, DeepONets, and FNOs, particularly when testing models on finer-scale data that do not lie in the training distribution. Similar trends were observed for the time-dependent diffusion problem and the advection-diffusion problem.
@@
A key distinguishing feature of NGOs is that, in addition to predicting the solutions of parametric PDEs, they return explicit representations of the parametric Green’s functions of such PDEs. We demonstrated that an explicit representation of the Green’s function enables the induction of biases in the network, such as preconditioning the output of the system network and the incorporation of conservation and dissipation laws, leading to improved accuracy of the predicted solutions. The efficacy of such inductive biases was demonstrated in the time-dependent diffusion problem, which allowed us to embed operator stability ensuring long-term autoregressive rollout and thereby pointwise accuracy in time, even when trained on a single time step. Furthermore, we demonstrated that Green’s functions inferred by NGOs can be used to construct effective matrix preconditioners for accelerating iterative solvers of large-scale problems.
@@
Although we provide a proof of concept for the use of NGOs as solution operators for linear PDEs, several research directions emerge from this work. One research direction includes enriching the space of parameters that are encoded in NGOs to solve PDEs in more general settings. Such settings can include varying geometries of the problem domain, as well as varying the type of boundary condition applied on the boundary segments. Another direction of research involves the customized use of NGOs to accelerate other algorithms. Of particular interest is the iterative use of NGOs within a fixed-point algorithm, such as Newton, to solve nonlinear problems, as well as the extension of the matrix preconditioning strategies for accelerating iterative solvers and time integration strategies for accelerating transient solvers. Lastly, we are also interested in exploring training strategies that would lead to even more accurate models.
@@ C.1 Steady Diffusion
Figure 22 shows the architectures of the DeepONet, FNO, and VarMiON. Figure 23 shows the architectures of the NGOs. The architectures of the system networks for the VarMiON and NGOs are shown in Figure 24. The U-Net architectures for the data NGO and VarMiON are identical with exception of the number of channels in the hidden layers. This is due to the extra $144\times 144$ linear layer in the VarMiON, which means that the U-Net must have fewer parameters in order to keep the total parameter count approximately the same.
Figure 22: From top to bottom: architectures of the DeepONet, FNO, and VarMiON.
@@
Table 10: Specific settings used for the different models of Section 4.3.2 regarding basis, quadrature and system net. $N_{x}$, $N_{y}$, $Q_{x}$ and $Q_{y}$ are the number of basis functions and quadrature points, respectively, in directions $x$ and $y$, respectively. The U-Nets and FNO follow the structure of Figures 22 and 24.
@@
  • CNN (convolutional neural network): the CNN follows the structure of the bottom model of Figure 24, without skip connections. The kernel, input and output shapes are mentioned in Tables 9, 10 and $1$.
  • U-Net: the U-Net follows the structure of the bottom model of Figure 24. The kernel, input and output shapes are mentioned in Tables 9, 10 and $1$.
  • FNO (fourier neural operator): the FNO follows the structure shown in Figure 22. The kernel, input and output shapes are mentioned in Tables 9, 10 and $1$.
@@ C.2 Time-Dependent Diffusion
Table 12: Specific settings used for the different models of Section 5.1 regarding basis, quadrature and system net. $N_{t}$, $N_{x}$, $N_{y}$, $Q_{t}$, $Q_{x}$ and $Q_{y}$ are the number of basis functions and quadrature points, respectively, in directions $t$, $x$ and $y$, respectively. The U-Nets and FNO follow the structure of Figures 22 and 24.
@@ Appendix G Relation to CNOs
Figures 26 and 27 show the encoding and decoding process as done in CNOs. Figure 28 shows the function basis corresponding to these encoding and decoding processes. This shows that CNOs use a specific function basis, which can actually be changed based on the problem. For example, while CNOs assume periodic domains, their Dirichlet kernel basis can be replaced by another basis, such as a B-spline basis, that conforms to the topology of the domain.

### v5

@@ 5.2 Advection-Diffusion
The next test case we consider is the advection-diffusion equation with inhomogeneous Dirichlet and Neumann boundary conditions:
@@
where $\left(\mathbf{c}\cdot\mathbf{n}\right)_{-}=\min\left(\mathbf{c}\cdot\mathbf{n},0\right)$. Compared to the pure diffusion equation, the advection-diffusion equation presents a more challenging problem due to a number of factors. First, the solution now depends nonlinearly on both the diffusion $\theta$ and the velocity field $\mathbf{c}$. Second, for large Péclet numbers (the ratio between the advection $\mathbf{c}$ and the diffusion $\theta$) the true solution will feature boundary layers near the Dirichlet boundaries, which may be difficult to resolve. Note that the boundary condition $\theta\nabla u\cdot\mathbf{n}-\left(\mathbf{c}\cdot\mathbf{n}\right)_{-}u=\eta$ differs from the Neumann boundary condition on the inflow part of the top and bottom boundaries (i.e. where $\mathbf{c}\cdot\mathbf{n}<0$). This change was made to facilitate stable finite element discretizations.
@@ 5.2.2 Results
Figure 20 shows the true solution alongside the solution predictions made by the different model architectures for one problem from the training data, and one from the testing data.
@@
Figure 20: The true solution and model predictions to one problem from the training data set (Figure 20(a)) and one problem from the testing data set (Figure 20(b)).
@@
To illustrate the scaling of the error with the data-set size we show in Figure 21 the effect that the amount of training data has on the model accuracy for FNOs, CNOs, data NGOs, and model NGOs. As the plot shows, both NGO models outperform the other models for each tested data set size, although the difference becomes small for large data sets. Figure 21 also shows the generalization gap, i.e. the difference between the training and testing error, which shows that NGOs consistently generalize better than FNOs, CNOs, and U-Nets. The VarMiON generalizes approximately as well as NGOs for relatively large data sets, but shows similar generalization to other models when the training data set is small.
@@
5.3 Nonlinear Diffusion
The final test case we consider is the nonlinear diffusion problem
$\begin{cases}\hfill-\nabla\cdot\left(\theta\left[u\right]\nabla u\right)=f&\text{on }\Omega=(0,1)^{2},\\ \hfill\theta\left[u\right]\nabla u\cdot\mathbf{n}=\eta&\text{on }\Gamma_{\mathrm{N}}=(0,1)\times\left\{0,1\right\},\\ \hfill\theta\left[u\right]u=g&\text{on }\Gamma_{\mathrm{D}}=\left\{0,1\right\}\times(0,1),\end{cases}$ (5.28)
where
$\theta\left[u\right](\mathbf{x})=\theta_{0}(\mathbf{x})+\alpha u(\mathbf{x})$ (5.29)
is a diffusion coefficient that depends linearly on the solution $u$. Here, $\theta_{0}(\mathbf{x})$ is a spatially varying, solution-independent part of the diffusion coefficient $\theta$, and $\alpha=0.1$ is a constant. Like in the steady diffusion problem 4, the domain $\Omega$ is the unit square, shown in Figure 4, Neumann boundary conditions $\eta(\mathbf{x})$ are applied on the top and bottom boundaries, and Dirichlet boundary conditions $g(\mathbf{x})$ are applied on the left and right boundaries.
Classical solution strategies to solve nonlinear PDEs like (5.28) often rely on iterative methods that formulate the solution to the nonlinear problem as a sequence of successive approximations to an associated linear parametric subproblem [46]. NGOs are well suited to accelerate such methods by directly replacing the expensive step of computing the solution to the parametrized linear subproblem within an iterative loop with an operator evaluation. To demonstrate the effectiveness of embedding NGOs into nonlinear iterative solvers, we consider the Picard iteration method [28, 46] for approximating solutions of (5.28). Starting from an initial guess $u^{(0)}=0$, the Picard iterates $\{u^{(k)}\}_{k\geq 1}$, where $k\in\{1,2,\ldots\}$, are defined by solving:
$\begin{cases}\hfill-\nabla\cdot\bigl(\theta[u^{(k-1)}]\nabla u^{(k)}\bigr)=f&\text{in }\Omega=(0,1)^{2},\\[4.0pt] \hfill\theta[u^{(k-1)}]\nabla u^{(k)}\cdot\mathbf{n}=\eta&\text{on }\Gamma_{\mathrm{N}}=(0,1)\times\{0,1\},\\[4.0pt] \hfill\theta[u^{(k-1)}]\,u^{(k)}=g&\text{on }\Gamma_{\mathrm{D}}=\{0,1\}\times(0,1).\end{cases}$ (5.30)
In this framework, a sequence of solutions $u^{(k)}$ of the linearized subproblems provides successive approximations of the solution $u$ of nonlinear PDE, given the previous iterate $u^{(k-1)}$.
Note that for each $k\in\{1,2,\ldots\}$, the system (5.30) is an instance of the parametric steady diffusion problem (4.1), obtained by setting $u\to u^{(k)}$ and $\theta\to\theta[u^{(k-1)}]$, and therefore the solution to (5.30) can be written in terms of a parametric Green’s operator as shown in (4.2). This observation enables the direct reuse of an NGO trained for the parametric linear problem, described in section 4.3.2, as a surrogate solver within each Picard iteration. However, we observed in our FEM tests for this problem that stabilization of our weak form, as described in remark 5, is necessary for the convergence of the Picard iterations. While the data NGOs pretrained on the linear problem described in 4.3.2 can be reused directly, the preconditioned data-free and model NGOs needed to account for such stabilization by adding the Nitsche stabilization terms (4.8) and (4.9) to the input matrix $\mathbf{F}$, the Neumann preconditioner $\mathbf{F}_{0}$ (described in section 4.2), and to the right-hand-side vector $\mathbf{d}$, respectively.
5.3.1 Results
In this section, we assess the effectiveness of embedding NGOs, trained on solutions of the parametric linear diffusion problem described in Section 4, within a Picard iteration (5.30) to solve the nonlinear diffusion problem (5.28) parametrized by $(\theta_{0},f,\eta,g)$. We compare the resulting performance against NO architectures from the literature that are capable of directly approximating solutions to nonlinear PDEs without resorting to iterative schemes. Specifically, we consider a U-Net (denoted NN), a DeepONet, and a FNO. For completeness, we also evaluate the performance of the U-Nets, DeepONets, and FNOs under two training and deployment strategies:
  • (i)
    Direct approximation of the solution operator for the parametric nonlinear problem. For this strategy the architectures and training procedure of the NNs, FNOs, DeepONets, are also identical to those in Section 4.3.2 (Appendices C.1.2 and D.1.2, respectively), however, these have been trained on nonlinear diffusion dataset F, described in Appendix B.4.
  • (ii)
    Training on the corresponding linear parametric problem and embedding the resulting models within Picard iterations, analogous to our approach with NGOs. For this strategy we reuse the trained models presented in Section 4.3.2 for the linear diffusion problem: their architectures, training procedure and dataset (dataset C) are highlighted in Appendices C.1.2, D.1.2 and B.1, respectively.
This comparison allows us to isolate the benefits of NGOs from those attributable solely to iterative deployment.
Figure 22 reports the relative error of the different network architectures as a function of the Picard iteration index $k$. These curves are overlaid with horizontal lines indicating the errors achieved by networks that directly approximate the solution operator of the nonlinear problem.
Figure 22: Relative $L^{2}$ test error over fixed point iteration number $k$ on (a) in distribution data with length scale $\lambda/L=1$ and (b) out of distribution data with $\lambda/L=0.1$ (where we used $L=1$), for a standalone NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO (an FNO is also added for comparison). For the NN, DeepONet and FNO, results are included for a direct as well as for an iterative solution approach. All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO) use the same 10 $\times$ 10 cubic B-spline basis, and are thus limited by the same $L^{2}$ projection error lower bound, indicated in black. Points and lines are averages, and error bars and shaded areas are 95% confidence intervals, over datasets of 1000 manufactured solutions generated in the same way as Dataset F (Appendix B.4).
In Figure 22(a) we present the results of in-distribution testing for a length scale $\lambda/L=1$. We observe that all networks deployed within Picard iterations initially mirror the convergence trajectory of the stabilized FEM; however, they eventually stagnate at higher errors. Among all models considered, the model NGO most closely tracks the convergence behavior of the stabilized FEM and achieves the lowest final error, followed by the data-free NGO and the data NGO. Furthermore, we also observe that the iterative use of the NN, FNO, and DeepONet converges to error levels comparable to those obtained by their direct nonlinear counterparts, indicating that iterative deployment does not confer a clear advantage over direct approximation strategy for these architectures in our tests.
Figure 22(b) reports the results of out-of-distribution testing for a finer length scale $\lambda/L=0.1$. Here we observe that all NGO variants converge in a single Picard iteration, with the errors of the data-free and model NGOs visually overlapping with the projection error. For clarity, Figure 22(b) omits the results for the NN, FNO, DeepONet, and VarMiON, as their relative errors are significantly larger. This trend is further corroborated in Figure 23, which summarizes the approximation errors of all considered neural operators across a range of test length scales $\lambda/L$.
Figure 23: Relative $L^{2}$ test error versus test data length scale $\lambda/L$ (where we used $L=1$) for a standalone NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO. All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO, also added as reference) use the same $10\times 10$ cubic B-spline basis, and are thus limited by the same $L^{2}$ projection error lower bound, indicated in black. Points and error bars are, respectively, averages and 95% confidence intervals on datasets of 1000 manufactured nonlinear diffusion solutions, generated in the same way as dataset F (Appendix B.4).
The results presented in Figure 23 closely mirror the trends observed for the linear diffusion problem in Figure 8. Specifically, NGOs maintain their superior performance when applied to nonlinear diffusion, exhibiting both higher in-distribution accuracy and significantly more robust generalization to fine-scale, out-of-distribution data. However, as NGOs function as structural components of an iterative solver rather than end-to-end mappings, they inherit the requirements of conventional numerical methods: their convergence relies on the availability of a suitable initial guess, and they require multiple operator evaluations per solution. This stands in contrast to other NO architectures, which provide a direct, albeit less generalizable, single-pass approximation.
Remark 9.
We observe that, for out-of-distribution fine-scale test data, the performance of the data-free NGO exhibits variability: among multiple independently and equivalently trained models, stable convergence was not consistently achieved. Consequently, we report results for the best-performing model selected from five such training runs. Developing inductive biases that explicitly promote stability of fixed-point iterations, analogous to the inductive biases for time-stepping stability introduced in Section 5.1.2, is an important direction for future work.
@@ 6 Conclusion
In this work, we have introduced the Neural Green’s Operator, a machine learning architecture for learning the solution operator to parametric families of linear PDEs. Our construction of NGOs is based on a finite-dimensional representation of Green’s functions. Such a representation enables the preservation of the linear action of Green’s operators on the inhomogeneity fields, while approximating the nonlinear dependence of the Green’s function on the coefficients of the PDE using neural networks that take weighted averages of such coefficients as input. This construction reduces the complexity of the problem from learning the entire solution operator and its dependence on all parameters to only learning the Green’s function and its dependence on the PDE coefficients. Moreover, taking weighted averages, rather than point samples, of input functions decouples the network size from the number of sampling points, enabling efficient resolution of multiple scales in the input fields. We distinguished three use cases of NGOs: the model NGO, which is used when the underlying PDE is known and training data is available; the data-free NGO, which is used when the underlying PDE is known, but training data is unavailable; and the data NGO, which is used when training data is available, but the model that underlies the data is unknown. For the steady diffusion problem, the accuracy of NGOs exceeds that of VarMiONs, DeepONets, and FNOs, particularly when testing models on finer-scale data that do not lie in the training distribution. Similar trends were observed for the time-dependent diffusion problem, the advection-diffusion problem and the nonlinear diffusion problem.
@@
We show that an explicit representation of the Green’s function enables the incorporation of inductive biases directly into the network architecture. In particular, preconditioning the output of the system network leads to improved solution accuracy, while for time-dependent problems, embedding operator stability together with conservation and dissipation laws enables stable long-term autoregressive rollouts and, consequently, pointwise accuracy in time. Furthermore, we demonstrated that Green’s functions inferred by NGOs can be used to construct effective matrix preconditioners for accelerating iterative solvers of large-scale problems. Preconditioners based on NGOs mimic popular two-level preconditioning strategies, making them applicable to a broad range of equations and linear solvers. Furthermore, NGO-based preconditioners were found to be significantly more effective than a similar strategy using DeepONets.
As a demonstration of the extension of NGOs to parametric nonlinear problems, we show that NGOs trained exclusively on solutions of associated linear PDEs can be effectively embedded within iterative nonlinear solvers. When supplied with a suitable initial guess, this approach yields accurate approximations to solutions of nonlinear parametric PDEs without retraining the model on nonlinear data. Moreover, the accuracy trends observed for the steady linear diffusion problem carry over to the nonlinear setting: NGOs consistently outperform VarMiONs, DeepONets, and FNOs, particularly in out-of-distribution regimes involving finer spatial scales than those seen during training. These results underscore the potential of NGOs to serve as reusable, modular components within classical numerical algorithms, effectively bridging linear training and nonlinear deployment.
@@
While we provide a proof of concept for the use of NGOs as solution operators for PDEs, several research directions emerge from this work. One research direction includes enriching the space of parameters that are encoded in NGOs to solve PDEs in more general settings. Such settings can include varying geometries of the problem domain, as well as varying the type of boundary condition applied on the boundary segments. Another direction of research involves the customized use of NGOs to accelerate other algorithms. Of particular interest is the extension of the matrix preconditioning strategies for accelerating iterative solvers for large-scale nonlinear systems, and time integration strategies for accelerating transient solvers. Lastly, we are also interested in exploring training strategies, such as multi-stage approaches [59], that would lead to even more accurate models.
@@
Acknowledgments
This project has received funding from the Eindhoven Artificial Intelligence Institute under the Exploratory Multidisciplinary AI Research program for the NGO-PDE project. The authors thank Barry Koren of Eindhoven University of Technology for his valuable insights and helpful comments.
@@ B.3 Advection-Diffusion
B.4 nonlinear Diffusion
Dataset F
Dataset F consists of 10000 manufactured GRF solutions, defined as in Equation B.1. We manufacture $\theta_{0}$ and $u$ by drawing the numbers $b$, $c$ and $\lambda$ from the distributions indicated in Table 9 and substituting them in Equation B.1. $\theta[u]$ is obtained by substitution of the manufactured $\theta_{0}$ and $u$ into Equation 5.29, where $\alpha=0.1$. The corresponding $f$, $\eta$ and $g$ were obtained by passing the manufactured $\theta[u]$ and $u$ through the respective PDE and boundary operators of 5.28.
Table 9: Distributions used for the Gaussian random field offset $b$, scaling $c$ and length scale $\lambda/L$ for the solution $u$ and material parameter $\theta_{0}$ for dataset C, manufactured using Equation B.1. $U$ is the uniform distribution.
Offset $b$ Scaling $c$ Length scale $\lambda/L$
Material parameter $\theta_{0}$ 1 $U(0,0.1)$ $U(0.5,1)$
Solution $u$ U(1,1)U(-1,1) $U(0,1)$ $U(0.5,1)$
@@ C.1 Steady Diffusion
Figure 24 shows the architectures of the DeepONet/RINO, FNO, and VarMiON. Figure 25 shows the architectures of the NGOs. The architectures of the system networks for the VarMiON and NGOs are shown in Figure 26. The U-Net architectures for the data NGO and VarMiON are identical with exception of the number of channels in the hidden layers. This is due to the extra $144\times 144$ linear layer in the VarMiON, which means that the U-Net must have fewer parameters in order to keep the total parameter count approximately the same.
Figure 24: From top to bottom: architectures of the DeepONet/RINO, FNO, and VarMiON.
@@
Table 11: Specific settings used for the different models of Section 4.3.2 regarding basis, quadrature and system net. $N_{x}$, $N_{y}$, $Q_{x}$ and $Q_{y}$ are the number of basis functions and quadrature points, respectively, in directions $x$ and $y$, respectively. The U-Nets and FNO follow the structure of Figures 24 and 26.
@@
  • CNN (convolutional neural network): the CNN follows the structure of the bottom model of Figure 26, without skip connections. The kernel, input and output shapes are mentioned in Tables 10, 11 and 12.
  • U-Net: the U-Net follows the structure of the bottom model of Figure 26. The kernel, input and output shapes are mentioned in Tables 10, 11 and 12.
  • FNO (fourier neural operator): the FNO follows the structure shown in Figure 24. The kernel, input and output shapes are mentioned in Tables 10, 11 and 12.
@@ C.2 Time-Dependent Diffusion
Table 13: Specific settings used for the different models of Section 5.1 regarding basis, quadrature and system net. $N_{t}$, $N_{x}$, $N_{y}$, $Q_{t}$, $Q_{x}$ and $Q_{y}$ are the number of basis functions and quadrature points, respectively, in directions $t$, $x$ and $y$, respectively. The U-Nets and FNO follow the structure of Figures 24 and 26.
@@ Appendix G Relation to CNOs
Figures 28 and 29 show the encoding and decoding process as done in CNOs. Figure 30 shows the function basis corresponding to these encoding and decoding processes. This shows that CNOs use a specific function basis, which can actually be changed based on the problem. For example, while CNOs assume periodic domains, their Dirichlet kernel basis can be replaced by another basis, such as a B-spline basis, that conforms to the topology of the domain.